Buildup each fraction or whole number so that it is equivalent to the fraction with the indicated denominator
5/7=?/98

Answer by fractalier(2101)

From
5/7 = ?/98
we ask, "What number did I multiply 7 by to get to 98?" Well, 14.
Thus you must multiply the 5 by 14 as well, so that
5/7 = 70/98